# A several variables Kowalski-Słodkowski theorem for topological spaces

### v1

@@ 4. A Multiplicative GKZ Theorem for Hardy spaces
Let $\mathcal{A}$ be a Banach algebra and $\phi:\mathcal{A}\to\mathbb{C}$ be a map satisfying
@@
If we restrict $F$ to $H^{\infty}(\mathbb{D})$, we can see that $F$ satisfies the hypotheses given in Theorem 4.1. As a result there exists a character $\psi_{F}$ on $H^{\infty}(\mathbb{D})$ such that $F(f)=\psi_{F}(f)$ for every invertible $f$ in $H^{\infty}(\mathbb{D})$. Let $f(z)=z+2$. Then $f$ is an invertible function in $H^{\infty}(\mathbb{D});$ therefore $F(f)=\psi_{F}(f)=\psi_{F}(z)+2.$ Also, by the hypotheses, $F(f)=w+2$ for some $w\in\mathbb{D}$. Then $\psi_{F}(z)=w\in\mathbb{D}$.
We first show that $\psi_{F}(f)=f(w)$ for all $f\in H^{\infty}(\mathbb{D})$. Let $f\in H^{\infty}(\mathbb{D})$. Then the function $g(z):=\big{(}f(z)-f(w)\big{)}/(z-w)$ is in $H^{\infty}(\mathbb{D})$ and $f-f(w)=(z-w)g$. By applying $\psi_{F}$ to the last equation, we get $\psi_{F}\big{(}f-f(w)\big{)}=\psi_{F}(z-w)\psi_{F}(g)=0$. Thus, $\psi_{F}(f)=f(w)$.
Let $f\in H^{p}(\mathbb{D})$ with no zeros in $\mathbb{D}$. For $10<r<$ 1, let us define $f_{r}(z)=f(rz),\ z\in\mathbb{D}$. Then each $f_{r}$ is in $H^{p}(\mathbb{D})$ and $f_{r}\to f$ in $H^{p}(\mathbb{D})$ as $r\to 1.$ Further, since $f$ does not have a zero in $\mathbb{D}$, each $f_{r}$ is invertible in $H^{\infty}(\mathbb{D}).$ Thus, $F(f_{r})=\psi_{F}(f_{r})=f_{r}(w)$ for every $r.$ Lastly, the continuity of $F$ yields $F(f)=\lim\limits_{r\to{1^{-}}}F(f_{r})=\lim\limits_{r\to{1^{-}}}f_{r}(w)=f(w).$ This completes the proof. ∎
@@ Acknowledgements
We thank the Mathematical Sciences Foundation, Delhi, for the support and facilities needed to complete the present work. The second author also thanks Shiv Nadar Institution of Eminence for partially supporting this research.

### v3

@@ 4. A Multiplicative GKZ Theorem for Hardy spaces
Let $\mathcal{A}$ be a unital Banach algebra and $\phi:\mathcal{A}\to\mathbb{C}$ be a map satisfying
@@
We consider the restriction of the map $F$ to the disc algebra $A(\mathbb{D})$ (a subset of $H^{p}(\mathbb{D})$ ). The restriction of the map $F$ to $A(\mathbb{D})$ (equipped with the sup-norm) satisfies the hypotheses of Theorem 4.1. As a result, there exists a character $\psi_{F}$ on $A(\mathbb{D})$ such that $F(f)=\psi_{F}(f)$ for every $f\in G_{1}(A(\mathbb{D})),$ the connected component of the set of invertible functions in $A(\mathbb{D})$ containing the constant function $1$. But, $G_{1}(A(\mathbb{D}))$ equals the set of all invertible functions in $A(\mathbb{D})$. Indeed, a simple adaptation of the arguments used in [9, Lemma 3.5.14] can be used to establish this fact. Therefore, we conclude that $F(f)=\psi_{F}(f)$ for each invertible function in $A(\mathbb{D}).$
Furthermore, the characters on the disc algebra are point evaluations, which implies that there exists a point $w\in\overline{\mathbb{D}}$ such that $F(f)=\psi_{F}(f)=f(w)$ for every invertible $f\in A(\mathbb{D}).$ Now, $g(z)=z+2$ is an invertible function in $A(\mathbb{D});$ therefore, $F(g)=\psi_{F}(g)=w+2.$ On the other hand, by the hypotheses, $F(g)=a+2$ for some $a\in\mathbb{D}$. Thus, we conclude that $w=a\in\mathbb{D}.$
Finally, let $f\in H^{p}(\mathbb{D})$ with no zero in $\mathbb{D}$. For $10<r<$ 1, let us define $f_{r}(z)=f(rz),\ z\in\overline{\mathbb{D}}$. Then each $f_{r}$ is in $H^{p}(\mathbb{D})$ and $f_{r}\to f$ in $H^{p}(\mathbb{D})$ as $r\to 1.$ Since $f$ is analytic on $\mathbb{D}$ and has no zero in $\mathbb{D}$, each $f_{r}$ is an invertible function in $A(\mathbb{D})$. Hence, $F(f_{r})=\psi_{F}(f_{r})=f_{r}(w)$ for every $r.$ Lastly, the continuity of $F$ yields $F(f)=\lim\limits_{r\to{1^{-}}}F(f_{r})=\lim\limits_{r\to{1^{-}}}f_{r}(w)=f(w).$ This completes the proof. ∎
@@ Acknowledgements
We thank the Mathematical Sciences Foundation, Delhi, for the support and facilities needed to complete the present work. The first named author thanks the University Grants Commission(UGC), India, for the support, and the second named author thanks Shiv Nadar Institution of Eminence for partially supporting this research.